OpenStudy (anonymous):

What is the discriminant of the given equation? x^2 - x - 2 = 0 A. 9 B. -7 C. 7 D. -9

OpenStudy (anonymous):

@gabbyalicorn do you know the answer? Because if not, i'm going to have to ask you to leave.

OpenStudy (anonymous):

lol there are 2 mlps in here

OpenStudy (gabbyalicorn):

You know what...I am trying to see if I can help you. -_- But looks like you're good.

OpenStudy (fluttershyk):

yeah that's what I was gonna say I love MLP FIM

OpenStudy (fluttershyk):

ok after long thought I have figured out the answer to your difficult problem

OpenStudy (anonymous):

None of the values as shown makes the sum zero. X does not satisfy any values.

OpenStudy (fluttershyk):

exactly the answer is -1

OpenStudy (anonymous):

-1 is the answer but it's not shown ...as one of the 4 answers...so there u go!

OpenStudy (anonymous):

None of you get the medal. You can't just give me an answer if it's not one of the answer choices.

OpenStudy (anonymous):

You must have calculated something wrong.

OpenStudy (fluttershyk):

I will explain it x2-x-2=0 (x-2)*(x+1)=0 x-2=0 x+1=0 x=-1 here is a graph